### v1

@@ Abstract
Accurate robot odometry is essential for autonomous navigation. While numerous techniques have been developed based on various sensor suites, odometry estimation using only radar and IMU remains an underexplored area. Radar proves particularly valuable in environments where traditional sensors, like cameras or LiDAR, may struggle, especially in low-light conditions or when faced with environmental challenges like fog, rain or smoke. However, despite its robustness, radar data is noisier and more prone to outliers, requiring specialized processing approaches. In this paper, we propose a graph-based optimization approach (https://github.com/robotics-upo/4D-Radar-Odom.git) using a sliding window for radar-based odometry, designed to maintain robust relationships between poses by forming a network of connections, while keeping computational costs fixed (specially beneficial in long trajectories). Additionally, we introduce an enhancement in the ego-velocity estimation specifically for ground vehicles, both holonomic and non-holonomic, which subsequently improves the direct odometry input required by the optimizer. Finally, we present a comparative study of our approach against existing algorithms, showing how our pure odometry approach inproves the state of art in most trajectories of the NTU4DRadLM dataset, achieving promising results when evaluating key performance metrics.
@@ 1 Introduction
Despite its advantages, radar presents several challenges that differentiate it from other sensors like LiDAR, with which it is commonly compared due to the similarity in data format—point clouds[9]. Radar data is inherently noisier and sparser than LiDAR data, leading to point clouds with a greater amount of irrelevant or incorrect information that complicates precise environmental interpretation. Its lower spatial and angular resolution also limits radar’s ability to capture fine structural details, making feature extraction techniques more challenging. Additionally, radar data is affected by artifacts such as specular reflections, multipath effects, and signal attenuation, resulting in both false detections and missed detections. Specular reflections occur when radar waves bounce off reflective surfaces at specific angles, causing the signal to deviate from a direct return path. This deviation can lead to false detections, where objects appear in incorrect positions, or missed detections when the signal does not return to the radar, hiding the actual object’s location. Multipath effects arise when signals reflect off multiple surfaces before returning to the radar, creating multiple echoes for the same object and distorting its perceived location and distance, with signal strength decreasing after each reflection. Signal attenuation also impacts radar reliability, as certain materials can absorb radar waves rather than reflecting them, resulting in additional missed detections. Together, these artifacts contribute to flickering between adjacent frames and ghost reflections, resulting in an unstable and potentially misleading perception of the environment. [7, 8].
@@
These challenges prevent the direct application of LiDAR-based techniques to radar data. Therefore, specialized algorithms are required to process radar’s noisier, lower-resolution data while mitigating artifacts that compromise data stability.
@@
In this paper, we present an odometry system that enables effective vehicle navigation in challenging environments using radar sensors and Inertial Measurement Units (IMUs). The system mitigates the effects of inherent sensor noise through a novel methodology for estimating the vehicle velocity via the Doppler effect, followed by a robust odometry estimation framework based on Pose Graph optimization. This robustness is achieved by forming interconnections among poses in a mesh-like structure, which reduces the impact of sensor noise.
1.1 Related work
@@
Additionally, sensor fusion is commonly employed across these methods to enhance performance by leveraging the strengths of different sensors. By combining radar data with other sensor modalities like cameras, IMUs, or thermal cameras, robustness is enhanced by compensating for the noise and sparsity of radar data with complementary information. For example, integrating inertial measurements from an IMU can improve the estimation of orientation and motion, aiding in the overall accuracy of the SLAM system. In [8], radar data is fused with thermal camera and IMU data, resulting in more reliable and accurate SLAM solutions in complex environments. Here, we focus on the loose combination of IMU and radar, that is, a solution that integrates IMU and radar data without the need for a strict interdependence in the data fusion process, allowing each sensor to contribute independently to the vehicle’s odometry estimation.
@@
1.2 Paper contribution
We propose an odometry system that takes advantage of pose graph optimization to improve the robustness and accuracy of 3D odometry using radar and IMU data. Our main contributions are:
  • A methodology for ego-velocity estimation that achieves higher accuracy reducing the impact of radar noise on the velocity estimation, specifically minimizing vertical drift. This method incorporates motion constraints derived from the vehicle’s kinematics (e.g., lateral movement restrictions in ground vehicles) directly into the estimation model. With IMU orientation data integrated between successive samples, the approach enhances the reliability of the velocity estimation. Additionally, RANSAC filtering mitigates the impact of dynamic objects, further refining the accuracy.
  • An odometry system designed with pose graph optimization that fuses radar and IMU measurements within a sliding window of the most recent n keyframes. Within this framework, poses in the window are interconnected, forming a mesh that establishes relative and local relationships. This optimization approach refines the vehicle’s trajectory by incorporating constraints from both IMU data and the alignment of radar scans at each keyframe. By leveraging these interconnections, the system enhances robustness against radar noise, as weaker connections are compensated by stronger ones within the window, effectively smoothing the trajectory and correcting drift. This structure maintains a constant computational cost while ensuring an accurate SLAM outcome.
@@ 2 Method overview
The proposed odometry approach is shown in the flow diagram (Fig. 1), it is divided into two main stages. The first handles filtering, Doppler velocity estimation, and the preprocessing and calibration of data from the IMU and radar sensors. It applies filtering techniques to the sensor readings and estimates the vehicle’s ego-velocity using the Doppler effect at each radar point, taking into account the vehicle’s degrees of freedom and reducing noise by excluding non-physically feasible movements. Additionally, dynamic objects are identified and discarded, ensuring only static elements contribute to the final velocity estimation.
The second stage is dedicated to the optimization process. Here, the preprocessed radar and IMU data, along with the aggregated ego-velocity estimate, serve as inputs for initial pose estimation, graph construction, and subsequent optimization as detailed in Section 1.2. Once optimized, poses outside the sliding optimization window are excluded from further modification and appended to the final trajectory, maintaining computational efficiency as the window advances.
3 Velocity Estimation with Ground Robot Constraints
@@
In order to have an accurate robot’s velocity estimation from 4D-radar sensing, we need to account for those points that belong to moving objects. Otherwise, the velocity estimation will integrate such objects as noise. The usual approach is to build a robust estimator based on RANSAC [6], detecting points belonging to moving objects and removing them from the estimator.
Despite the averaging effect produced by solving the system of equations, the resulting velocity is usually noisy due to the nature of the sensor, affecting both the points and the Doppler speed, which subsequently affects the estimated velocity. However, when it comes to ground robots, the vehicle’s movement can be restricted by considering that it cannot have independent motion in the Z-axis component, with changes in altitude being limited by the vehicle’s tilt angle. Similarly, if the ground robot is non-holonomic, it can only move longitudinally, with the lateral component being dependent on the yaw and pitch rotations. Based on these considerations, an improvement in the estimation of velocity is proposed by incorporating movement restrictions for the vehicle.
@@ 3.1 Improved Velocity Estimation Using Orientation
To improve upon the method presented in [6], we introduce a refinement based on the vehicle’s previous and current orientations, thus obtaining an estimate of the vehicle’s change in orientation at the time of the scan. Considering that the vehicle is ground-based, the movement should occur primarily in the longitudinal $V_{x}$ and lateral $V_{y}$ axes (holonomic). Instead of estimating all components independently, we can relate them through this change in orientation. The velocity components $v_{x}$, $v_{y}$, and $v_{z}$ are therefore related as follows:
@@
where $\theta$ represents the pitch angle and $\phi$ the roll angle, both obtained from the IMU measurements. In this way, we reduce the system of equations by one degree of freedom and, more importantly, constrain the value of $\mathbf{v}$ so that it accounts for feasible robot velocities. Thus, if the vehicle is navigating over a relatively flat area, even though there may be noise in the measurement, the noise in $v_{z}$ can be addressed and partially removed. The system of equations to be solved becomes:
@@
On the other hand, non-holonomic vehicles, such as cars, can only move in their longitudinal direction. In this case, the system of equations is significantly constrained, as shown in (5).
@@
Here, spherical coordinates are used to relate the components of the velocity vector, with $\psi$ and $\theta$ as the yaw and pitch rotations from the IMU. This also reduces the number of unknowns to just one: the magnitude of the velocity $V$.
@@ 4.2 GICP-based Alignment of RADAR point-clouds
$w=\frac{1.0}{2f+\epsilon}$ (9)
where $f$ represents the GICP fitness score and $\epsilon$ is a small constant (e.g., $10^{-6}$) to prevent division by zero.
This weight is applied to the GICP constraint, thereby attenuating the influence of constraints with higher fitness scores during the optimization process. Additionally, alignments that yield fitness scores exceeding a predefined threshold are discarded, as they typically do not contribute to improving the initial pose estimation. Although such constraints receive a reduced weight via the factor $w$, they are excluded to prevent potential degradation of the optimization outcome.
To enhance robustness, the SLAM algorithm incorporates a strategy wherein poses are interconnected through multiple alignments within the sliding window (see Fig. 2). If a particular alignment fails to converge or produces a high fitness score, the associated constraint’s weight is diminished, thereby minimizing its impact on trajectory optimization. However, due to the interconnected nature of poses within the mesh of alignments, successful alignments from other keyframes maintain the integrity of the optimization process.
@@ 4.3 Graph Construction and Optimization
The optimization graph is constructed by defining a state vector $\mathbf{X}$ that includes the individual position and orientation parameters of each pose in the global system refernce within a window of the last $n$ poses, as follows:
@@
Where each pose $\mathbf{x}_{i}$ in the window is defined as $\mathbf{x}_{i}=[x_{i},y_{i},z_{i},roll_{i},pitch_{i},yaw_{i}]$. These parameters are optimized by minimizing cost functions associated with Generalized ICP (GICP) constraints. Two primary cost functions derived from the GICP constraint are applied, where each function enforces alignment between successive poses in the window, thus refining both position and orientation estimates across the trajectory.
@@
    $r_{\text{pos}}=w_{p}\cdot w\cdot\left(t_{\text{GICP}}-q_{a}^{o^{-1}}\left(t_{b}^{o}-t_{a}^{o}\right)\right)$ (11)
    where $t_{a}^{o}$ and $t_{b}^{o}$ represent the translation vectors of poses $a$ and $b$, respectively. The weight $w$ is directly applied to the residual to modulate its influence based on the fitness score as shown in 4.2, ensuring that less reliable constraints yield a diminished impact on the final optimization solution.
  • Orientation Cost Function: This function calculates the residuals based on the discrepancy between the GICP-derived quaternion and the estimated quaternion. The rotational error $\ {r}_{\text{ori}}$ is given by:
    $\ {r}_{\text{ori}}=\ {w}_{o}\cdot w\cdot\left((\ {q}_{\text{GICP}}^{-1}\cdot\ {q}_{a}^{o^{-1}}\cdot\ {q}_{b}^{o})-q_{I}\right)$ (12)
    where $q_{a}^{o}$ and $q_{b}^{o}$ are the orientation quaternions of poses $a$ and $b$, respectively, and $q_{\text{GICP}}$ refers to the relative rotation estimated by the GICP algorithm between the two poses. Here, $q_{I}$ denotes the identity quaternion. Analogous to the position cost function, the weight $w$ is applied directly to the residual to adjust its impact in the optimization process, modulating its influence in the optimization to favor constraints with higher reliability.
@@
Additionally, the following two more constraints are incorporated into the optimization process:
  • IMU Pitch and Roll Constraint: This cost function enforces consistency for the pitch and roll angles provided by the IMU, remain consistent throughout the optimization. Since the IMU-derived pitch and roll angles typically exhibit bounded error, this constraint helps maintain the poses within an acceptable range, enhancing the stability of the pose estimates.
@@
    This constraint is applied locally to each keyframe within the optimization window and is weighted adaptively based on the reliability of the GICP-derived orientation constraint. Specifically, when the GICP constraint has a lower weight—indicating reduced reliability—the IMU constraint exerts a stronger influence to maintain pitch and roll consistency. In contrast, when the GICP constraint is deemed reliable (i.e., with a high fitness score), the IMU constraint’s weight is reduced, allowing for subtle adjustments to pitch and roll that improve alignment.
  • Vehicle Dynamics Constraint: To reflect the ground vehicle’s physical limitations, this constraint restricts vertical (Z-axis) movements inconsistent with the vehicle’s dynamics. This is particularly advantageous given that GICP often introduces noise into the z-component, occasionally resulting in unrealistic vertical displacements. By imposing this constraint, the optimization better aligns with the physical motion constraints of ground vehicles, reducing the likelihood of non-physical z-axis shifts.
@@
    where $t_{\text{odom}}$ denotes the z-axis translation between keyframes $a$ and $b$, referenced in the coordinate frame of pose $a$. This translation measurement is derived from direct odometry and limits excessive z-axis displacements that may arise from GICP-induced noise.
Ceres solver is employed to minimize the weighted residuals, constructing a graph where each keyframe within the optimization windows serves as a node with associated parameter blocks, such as orientation and translation. Each parameter block is linked to custom cost functions and kernels that handle specific constraints—such as those from GICP, IMU, and vehicle dynamics. These cost functions integrate fitness scores to adjust constraint weights dynamically, with the associated kernels mitigating the influence of outliers by down-weighting high-residual values. The optimization process uses a sparse normal Cholesky linear solver and a Levenberg-Marquardt trust region strategy, efficiently balancing computational cost and accuracy.
When GICP constraints are highly precise, they are prioritized, guiding the trajectory estimation with minimal correction from other sources. However, as the reliability of the GICP constraints decreases, the IMU and vehicle dynamics constraints assume a greater role, ensuring consistent and physically plausible pose adjustments. This adaptive approach enables the Ceres solver to deliver an accurate and robust vehicle trajectory estimation within a fixed computational budget.

### v2

@@ Abstract
Accurate robot odometry is essential for autonomous navigation. While numerous techniques have been developed based on various sensor suites, odometry estimation using only radar and IMU remains an underexplored area. Radar proves particularly valuable in environments where traditional sensors, like cameras or LiDAR, may struggle, especially in low-light conditions or when faced with environmental challenges like fog, rain or smoke. However, despite its robustness, radar data is noisier and more prone to outliers, requiring specialized processing approaches. In this paper, we propose a graph-based optimization approach (https://github.com/robotics-upo/4D-Radar-Odom.git) using a sliding window for radar-based odometry, designed to maintain robust relationships between poses by forming a network of connections, while keeping computational costs fixed (specially beneficial in long trajectories). Additionally, we introduce an enhancement in the ego-velocity estimation specifically for ground vehicles, both holonomic and non-holonomic, which subsequently improves the direct odometry input required by the optimizer. Finally, we present a comparative study of our approach against existing algorithms, showing how our pure odometry approach inproves the state of art in all trajectories of the NTU4DRadLM dataset, achieving promising results when evaluating key performance metrics.
@@
Published in Journal of Intelligent and Robotic Systems, Volume 111, Article 107 (2025).
DOI: 10.1007/s10846-025-02301-9.
@@ 1 Introduction
Despite its advantages, radar presents several challenges that differentiate it from other sensors like LiDAR, with which it is commonly compared due to the similarity in data format—point clouds[9]. Radar data is inherently noisier and sparser, leading to point clouds with more irrelevant or incorrect information, complicating precise environmental interpretation. Its lower spatial and angular resolution also limits radar’s ability to capture fine structural details, making feature extraction techniques more challenging. Additionally, radar data is affected by artifacts such as specular reflections, multipath effects, and signal attenuation, leading to false or missed detections. Specular reflections occur when radar waves bounce off reflective surfaces at specific angles, causing false detections where objects appear in incorrect positions or missed detections when the signal fails to return. Multipath effects arise when signals reflect off multiple surfaces before returning to the radar, creating multiple echoes for the same object and distorting its perceived location and distance, with signal strength decreasing after each reflection. Signal attenuation also impacts radar reliability, as certain materials can absorb radar waves rather than reflecting them, resulting in additional missed detections. These artifacts contribute to flickering between adjacent frames and ghost reflections, resulting in an unstable and potentially misleading perception of the environment. [7, 8].
@@
These challenges prevent the direct application of LiDAR-based techniques to radar data. Therefore, specialized algorithms are required to process radar’s noisier, lower-resolution data while mitigating artifacts that compromise data stability. One critical aspect affected by radar noise is ego-velocity estimation, particularly in the vertical component, where lower spatial resolution and multipath effects contribute to increased drift and instability. Accurately estimating velocity from radar measurements is essential for reliable vehicle localization and motion prediction.
@@
In this paper, we present an odometry system that enables effective vehicle navigation in challenging environments using radar sensor and an Inertial Measurement Unit (IMU). The system mitigates the effects of inherent sensor noise through a novel methodology for estimating the vehicle velocity via the Doppler effect, followed by a robust odometry estimation framework based on Pose Graph optimization. Therefore, our main contributions are:
  • A methodology for ego-velocity estimation that achieves higher accuracy by reducing the impact of radar noise on velocity estimation, specifically minimizing vertical drift. This method incorporates motion constraints derived from the vehicle’s kinematics (e.g., vertical movement restrictions in ground vehicles) directly into the estimation model. With IMU orientation data integrated between successive samples, the approach enhances the reliability of the velocity estimation. Additionally, RANSAC filtering mitigates the impact of dynamic objects, further refining accuracy.
  • An odometry system designed with pose graph optimization that fuses radar and IMU measurements within a sliding window of the most recent n keyframes. Within this framework, poses in the window are interconnected, forming a mesh that establishes relative and local relationships. This optimization approach refines the vehicle’s trajectory by incorporating constraints from both IMU data and the alignment of radar scans at each keyframe. By leveraging these interconnections, the system enhances robustness against radar noise, as weaker connections are compensated by stronger ones within the window, effectively smoothing the trajectory and correcting drift. This structure maintains a constant computational cost while ensuring an accurate enhanced odometry outcome.
2 Related work
@@
Additionally, sensor fusion is commonly employed across these methods to enhance performance by leveraging the strengths of different sensors. By combining radar data with other sensor modalities like cameras, IMUs, or thermal cameras, robustness is enhanced by compensating for the noise and sparsity of radar data with complementary information. For example, integrating inertial measurements from an IMU can improve the estimation of orientation and motion, aiding in the overall accuracy of the SLAM system. In [8], radar data is fused with thermal camera and IMU data, resulting in more reliable and accurate SLAM solutions in complex environments. Sensor fusion strategies are typically categorized as ”loosely coupled” or ”tightly coupled,” depending on how the sensor data is integrated. In a tightly coupled approach, raw measurements from different sensors (e.g., IMU acceleration, radar Doppler, or even GNSS signals) are directly processed together in a unified estimation framework, often using filters or optimization-based techniques. This allows for deeper integration but increases computational complexity and dependency on accurate sensor models. In contrast, a loosely coupled approach processes each sensor independently to estimate intermediate states, such as velocity or pose, which are later combined at a higher level. This modular structure makes the system more adaptable and computationally efficient, at the cost of potentially weaker integration of constraints. Here, we focus on the loose combination of IMU and radar, that is, a solution that integrates IMU and radar data without the need for a strict interdependence in the data fusion process, allowing each sensor to contribute independently to the vehicle’s odometry estimation. Specifically, radar provides velocity estimations via the Doppler effect, while IMU data is used to track orientation and motion dynamics, which are then combined in the odometry framework. This approach balances robustness and computational efficiency, making it well-suited for real-time applications in challenging environments.
@@ 3 Method overview
The proposed odometry approach is shown in the flow diagram (Fig. 1), it is divided into two main stages: preprocessing and velocity estimation, followed by optimization. The first handles the calibration of data from the IMU and radar sensors, preprocessing, Doppler velocity estimation and filtering. It applies filtering techniques to the sensor readings and estimates the vehicle’s ego-velocity using the Doppler effect at each radar point, taking into account the vehicle’s degrees of freedom and reducing noise by excluding non-physically feasible movements. This method robustly filters out outliers by detecting inconsistent motion patterns, ensuring that only static elements contribute to the final velocity estimation. Furthermore, the IMU is an essential component of the velocity estimation process, as its orientation measurements directly constrain the system of equations used to compute the vehicle’s motion. These constraints help to reduce uncertainty, particularly in the vertical component, and improve the overall robustness of the estimation.
The second stage focuses on the optimization process. The preprocessed radar and IMU data, along with the refined ego-velocity estimate, are used as input for initial pose estimation and graph construction. A pose graph is incrementally built, where radar scans and IMU readings define constraints that link consecutive keyframes. These constraints are formulated to enforce consistency in the estimated trajectory while accounting for the uncertainty inherent in radar data. The optimization framework then refines the vehicle’s trajectory by minimizing errors across the pose graph. To maintain computational efficiency, a sliding window approach is employed: Only the most recent n keyframes are included in the optimization, and once poses exit this window, they are considered fixed and appended to the final trajectory without further modification. This ensures that the system remains scalable while preserving global consistency.
4 Velocity Estimation with Ground Vehicle Constraints
@@
In order to have an accurate robot’s velocity estimation from 4D-radar sensing, we need to account for those points that belong to moving objects. Otherwise, the velocity estimation will integrate such objects as noise. The usual approach is to build a robust estimator based on RANSAC [6], detecting points belonging to moving objects and removing them from the estimator. Despite these filtering steps, Doppler-based velocity estimation can still suffer from vertical drift and inconsistencies due to sensor noise. To improve robustness, we incorporate motion constraints based on IMU measurements, as detailed in the next section.
@@ 4.1 Improved Velocity Estimation Using Orientation
The Doppler-based velocity estimation method provides an initial estimate by solving a system of linear equations derived from radar measurements. However, this estimation is susceptible to noise and inconsistencies, particularly in the vertical component. To mitigate these issues, we incorporate motion constraints based on the vehicle’s physical movement model, ensuring that only feasible velocities are considered. Ground robots exhibit structured motion patterns that can be leveraged to refine velocity estimation:
  • Their movement along the Z-axis is constrained by the vehicle’s pitch and roll angles, meaning vertical velocity variations are primarily influenced by changes in orientation.
  • Non-holonomic constraints restrict lateral motion, meaning that movement is largely constrained to the forward direction
Since these constraints depend on the vehicle’s orientation, the IMU plays a crucial role in providing real-time measurements. Using these properties, we reformulate the velocity estimation problem by incorporating orientation constraints. Additionally, as in the basic Doppler model, a RANSAC-based approach is applied, which identifies as outliers those points with velocity different from the majority. Since most points in the environment correspond to static elements with a velocity that remains within a consistent range relative to the vehicle, dynamic objects exhibit atypical velocities that deviate significantly from this pattern. These outliers are filtered out to improve the accuracy of the estimation. Additionally, RANSAC helps mitigate the impact of radioelectrical effects such as specular reflections and multipath effects, which can also introduce inconsistent Doppler measurements.
4.1.1 Orientation Constraints for Holonomic Vehicles
For holonomic ground robots, motion primarily occurs along the longitudinal $V_{x}$ and lateral $V_{y}$ axes. Instead of estimating all components independently, we can relate them through this change in orientation. The velocity components $v_{x}$, $v_{y}$, and $v_{z}$ are therefore related as follows:
@@ 4.1.1 Orientation Constraints for Holonomic Vehicles
where $\theta$ represents the pitch angle increment and $\phi$ the roll angle increment, both from the IMU measurements, corresponding to the orientation change between the current and previous radar scan. In this way, we reduce the system of equations by one degree of freedom and, more importantly, constrain the value of $\mathbf{v}$ so that it accounts for feasible robot velocities. Thus, if the vehicle is navigating over a relatively flat area, even though there may be noise in the measurement, the noise in $v_{z}$ can be addressed and partially removed. Therefore, the system of equations to be solved, shown in (2), with the corresponding modification in the velocity vector seen in (3), becomes:
@@
4.1.2 Orientation Constraints for Non-Holonomic Vehicles
For non-holonomic vehicles, such as cars, motion is posible only in their longitudinal direction. This means that the velocity vector $\mathbf{v}$, initially defined by it components $v_{x}$, $v_{y}$, and $v_{z}$, can be reduced to a single component, $v_{x}$. To estimate the velocity in this scenario, we replace $v_{x}$ with a vector of magnitude $V$, wich can be represented using spherical coordinates. The resulting system of equation is then significantly constrained as shown in (5).
@@ 4.1.2 Orientation Constraints for Non-Holonomic Vehicles
Here, $\psi$ corresponds to the yaw angle increment and $\theta$ to the pitch angle increment, both obtained from IMU measurements, corresponding to the orientation change between the current and previous radar scan. This also reduces the number of unknowns to just one: the magnitude of the velocity $V$.
@@ 5.2 GICP-based Alignment of RADAR point-clouds
${\color[rgb]{0,0,0}w_{\text{GICP}}}=\frac{1.0}{2f+\epsilon}$ (9)
where $w_{\text{GICP}}$ is the weight aplied to the corresponding cost function, $f$ represents the GICP fitness score and $\epsilon$ is a small constant (e.g., $10^{-6}$) to prevent division by zero.
This weight is applied to the GICP constraint, thereby attenuating the influence of constraints with higher fitness scores during the optimization process, while emphasizing those with lower fitness scores. Additionally, alignments that yield fitness scores exceeding a predefined threshold $f_{\text{th}}$ are discarded, as they typically do not contribute to improving the initial pose estimation. Although such constraints receive a reduced weight via the factor $w$, they are excluded to prevent potential degradation of the optimization outcome.
To enhance robustness, the odometry algorithm incorporates a strategy where poses are interconnected through multiple alignments within the sliding window (see Fig. 2). If a particular alignment fails to converge or produces a high fitness score, the associated constraint’s weight is diminished, thereby minimizing its impact on trajectory optimization. However, due to the interconnected nature of poses within the mesh of alignments, successful alignments from other keyframes maintain the integrity of the optimization process.
@@ 5.3 Graph Construction and Optimization
The pose graph optimization is formulated as a nonlinear least-squares problem, where the state vector $\mathbf{X}$ is estimated by minimizing a set of cost functions, which define the residuals by relating the poses to each other and to external information. The state vector consists of individual position and orientation parameters for each pose within a sliding window of the last n poses.
@@
where each pose $\mathbf{x}_{i}$ in the window is defined as $\mathbf{x}_{i}=[x_{i},y_{i},z_{i},q_{w_{i}},q_{x_{i}},q_{y_{i}},q_{z_{i}}]$. To handle the non-Euclidean nature of quaternion-based orientation representations, the optimization is performed on a manifold. Specifically, quaternion updates are carried out in the local tangent space to ensure minimal and valid perturbations during optimization, following common practice in odometry and state estimation problems.
The optimization process seeks to find the optimal state $\mathbf{X}^{*}$ by minimizing the weighted sum of residuals of each cost function:
$\mathbf{X}^{*}=\arg\min_{\mathbf{X}}\sum_{i}r_{i}^{2}$ (11)
where $r_{i}$ represents the residuals associated with different cost function.
To efficiently solve this problem, we employ the Ceres Solver [33], which optimizes the state vector by minimizing the cost functions within a sliding window encompassing the last n poses of the trajectory. Within this window, each pose is interconnected with the previous ones, forming an optimization graph, as shown in Figure 3, where each keyframe serves as a node with associated parameter blocks.
As the trajectory progresses, older poses are removed from the optimization graph while new ones are incorporated, ensuring a balance between global consistency and adaptability to new sensor data. This sliding-window approach allows recent observations to have a stronger influence on the estimated trajectory while still maintaining coherence with past constraints.
The optimizer is configured to use a Levenberg-Marquardt trust-region strategy combined with a sparse normal Cholesky solver to balance computational cost and accuracy. Additionally, a Tukey robust kernel is applied to mitigate the influence of outliers, particularly those arising from noisy sensor measurements. The optimization runs for a maximum of 100 iterations, with convergence criteria set to their default values (function_tolerance: $10^{-6}$, gradient_tolerance: $10^{-10}$).
Figure 3: Optimization graph. The highlighted region shows new constraints added with the latest pose, while the shaded area represents existing constraints. Colors indicate radar measurements (purple), IMU measurements (orange), optimized poses (blue), and raw odometry transformations (gray).
5.3.1 Cost Function Definitions
Two primary cost functions derived from the GICP constraint are applied, where each function enforces alignment between successive keyframe radar measurements in the window, thus refining both position and orientation estimates across the trajectory. As shown in Figure 3, this constraint, marked by a purple triangle, links radar measurements to provide a refined transformation between poses, as detailed in Section 5.2.
@@ 5.3.1 Cost Function Definitions
    $r_{\text{pos}}=w_{p}\cdot w_{\text{GICP}}\cdot\left(t_{\text{GICP}}-q_{a}^{o^{-1}}\left(t_{b}^{o}-t_{a}^{o}\right)\right)$ (12)
    where $t_{a}^{o}$ and $t_{b}^{o}$ represent the translation vectors of poses $a$ and $b$, respectively, and $q_{a}^{o}$ the global quaternion orientation of pose a. The weight $w_{\text{GICP}}$ is directly applied to the residual to modulate its influence based on the fitness score as shown in 5.2, ensuring that less reliable constraints yield a diminished impact on the final optimization solution.
  • Orientation Cost Function: This function quantifies the residuals by evaluating the discrepancy between the GICP-derived quaternion and the estimated quaternion. The rotational error ${r}_{\text{ori}}$ is given by:
    ${r}_{\text{ori}}=\ {w}_{o}\cdot w_{\text{GICP}}\cdot\left((\ {q}_{\text{GICP}}^{-1}\cdot\ {q}_{a}^{o^{-1}}\cdot\ {q}_{b}^{o})-q_{I}\right)$ (13)
    where $q_{a}^{o}$ and $q_{b}^{o}$ are the global orientation quaternions of poses $a$ and $b$, respectively, and $q_{\text{GICP}}$ refers to the relative rotation estimated by the GICP algorithm between the two poses. Here, $q_{I}$ denotes the identity quaternion, ensuring that the rotational residual is zero when no rotation is present. Analogous to the position cost function, the weight $w_{\text{GICP}}$ is applied directly to the residual to adjust its impact in the optimization process, modulating its influence in the optimization to favor constraints with higher reliability.
@@
Additionally, the following two constraints are also incorporated into the optimization process:
  • IMU Pitch and Roll Cost Function: This cost function, shown in Figure 3 as orange elements, enforces consistency for the pitch and roll angles provided by the IMU, remain consistent throughout the optimization. Since the IMU-derived pitch and roll angles typically exhibit bounded error, this constraint helps maintain the poses within an acceptable range, enhancing the stability of the pose estimates. These values correspond to the orientation estimates provided directly by the commercial IMUs, which typically includes onboard EKF filtering to fuse gyroscope, accelerometer, and magnetometer data. This process compensates for sensor biases and provides stable roll and pitch angles, usually with errors below $0.5^{\circ}$. In contrast, yaw, being more affected by magnetic perturbations, is handled indirectly through the alignment-based constraints in the graph.
@@
    This constraint is applied locally to each keyframe within the optimization window and is weighted adaptively based on the reliability of the GICP-derived orientation constraint. Specifically, when the GICP constraint has a lower weight, indicating reduced reliability, the IMU constraint exerts a stronger influence to maintain pitch and roll consistency. In contrast, when the GICP constraint is deemed reliable (i.e., with a high fitness score), the IMU constraint’s weight is reduced, allowing for subtle adjustments to pitch and roll that improve alignment.
  • Vehicle Dynamics Cost Function: To reflect the ground vehicle’s physical limitations, this constraint, shown in Figure 3 (gray elements), restricts vertical (Z-axis) movements inconsistent with the vehicle’s dynamics. This is particularly advantageous given that GICP often introduces noise into the z-component, occasionally resulting in unrealistic vertical displacements. By imposing this constraint, the optimization better aligns with the physical motion constraints of ground vehicles, reducing the likelihood of non-physical z-axis shifts.
@@
    where $t_{\text{odom}}$ denotes the z-axis translation between keyframes $a$ and $b$, referenced in the coordinate frame of pose $a$. This translation measurement is derived from direct odometry and limits excessive z-axis displacements that may arise from GICP-induced noise. This constraint incorporates the effect of the refined velocity estimation shown in Section 4 into the optimization process. Since the initial odometry guess is directly influenced by the velocity refinement, the enforced limitation on vertical displacement ensures that the optimization process remains consistent with the improved motion estimate. As a result, even when GICP introduces noise in the z-component, the optimization process remains consistent with the physically valid motion dictated by the velocity-based odometry, preventing unrealistic vertical shifts while maintaining consistency with the vehicle’s dynamics. This constraint could also be extended to non-holonomic ground vehicles by incorporating the lateral y-component into the residual. However, it has been applied specifically to the z-component to simplify the optimization problem, as GICP primarily intoduces errors in the vertical direction.
The solver prioritizes GICP constraints when they exhibit high reliability, ensuring that the trajectory estimation relies primarily on accurate geometric alignments. However, as the reliability of the GICP constraints decreases, the IMU and vehicle dynamics constraints assume a greater role, ensuring consistent and physically plausible pose adjustments. By dynamically adjusting constraint weights based on their reliability, the optimization balances accuracy and robustness, maintaining computational efficiency within a fixed budget.